# Hunting for Polluted White Dwarfs and Other Treasures with Gaia XP Spectra and Unsupervised Machine Learning

### v1

@@ 1 Introduction
The death of a low-mass ($\lesssim$ 8 M) main-sequence star culminates in the ejection of its outer layers in a planetary nebula and the collapse of its core into a white dwarf (WD). Typical mass WDs are extremely dense and have very high surface gravities (log $g$ $\sim$ 8.0, central density $\approx 10^{6}$ $\mathrm{g/cm^{3}}$), equivalent to about 100,000 times that of Earth. As such, WDs are chemically stratified, meaning lighter elements like H and He rise to the surface and heavier elements (e.g., C, O, Ca, Mg, Fe, etc.) sink to the core. The majority of WDs are expected to consist of carbon-oxygen cores and thin upper layers of He and H that make up only $\sim$ 1% of the total WD mass. About 70% of all WDs have H-dominated atmospheres (DA) and the other $\sim$ 15% are He-dominated (DB). We typically expect the presence of H or He absorption lines in a WD spectrum, or no spectral lines if the WD is cold enough to no longer excite atoms above their ground state ($\lesssim$ 11,000 K for DBs and $\lesssim$ 5,000 K for DAs). However, some WDs, especially cooler WDs, have been found with absorption features from heavier elements in their atmospheres. This has been interpreted as evidence that planets can survive the mass loss stage of their host stars (Debes & Sigurdsson, 2002; Jura, 2003; Zuckerman et al., 2007; Koester et al., 2014).
@@
About 25 – 50% of all WDs below 20,000 K are polluted by heavy metals(Zuckerman et al., 2010; Koester et al., 2014; Badenas-Agusti et al., 2023). Observations of intact planets around WDs (Vanderburg et al., 2020; Mullally et al., 2024) and transiting planetary debris (Vanderburg et al., 2015; Vanderbosch et al., 2020, 2021) further lends the notion that planets are the main culprits for the presence of heavy metals in WD atmospheres. Planets and planetesimals that survive the late-stage evolution of their host stars and enter the WD’s Roche radius, possibly through orbital disruptions during mass loss in the asymptotic giant phase and dynamical interactions with other planets (Maldonado et al., 2020), become tidally disrupted and the debris is subsequently accreted onto the WD. Since the discovery of the first polluted WD, nearly 1400 WDs have been spectrally classified as metal-polluted (DZ, or DAZ, DZA, DBZ, DZB, etc. depending on the predominant lines in their spectra) (Dufour et al., 2017), with a majority only exhibiting Ca pollution.
@@
The most recent Gaia data release (DR3) in 2022 (Gaia Collaboration et al., 2023) revealed nearly 1.5 billion sources with full astrometry, an increase of nearly 136 million sources from the prior data release in 2019. The plethora of data from Gaia has enabled advances in areas such as Galactic archaeology (Poggio et al., 2018; Lucey et al., 2023; Deason & Belokurov, 2024), stellar evolution (Jao et al., 2018; Fouesneau et al., 2023), and measuring cosmic distances (Bailer-Jones et al., 2021; Riess et al., 2021; Ripepi et al., 2023). Gaia has also been pivotal in the field of WD astrophysics, increasing the number of WDs by nearly 100,000 after the second data release (Gentile Fusillo et al., 2019), which contributed to such advancements as the discovery of the WD crystallization branch and cooling anomaly (Tremblay et al., 2019; Cheng et al., 2019) and the development of a technique for finding new variable WDs using Gaia excess photometric scatter (Guidry et al., 2021). The Gaia observations consist of broadband (G), blue passband (BP), and red passband (RP) photometry that cover the wavelength ranges 330–1050 nm, 330–680 nm, and 640–1050 nm, respectively. In earlier data releases, BP and RP were simple photometric magnitudes. However, Gaia DR3 integrated low resolution ($R=\lambda/\Delta\lambda\sim$ 70) blue and red grism spectra to create the BP and RP photometry, and released these spectra for 220 million sources. Despite the low resolution of the BP and RP (XP) spectra, they are invaluable for characterizing atmospheric parameters like effective temperature (Teff), surface gravity (log $g$), metallicity, and line identification for vast quantities of stars (Andrae et al., 2023; Zhang et al., 2023; Vincent et al., 2023).
@@ 3.1 Categorizing Gaia DR3 WDs with UMAP
It has already been shown that the Gaia XP spectra can be exploited through machine learning to label and infer stellar parameters for vast quantities of stars (Andrae et al., 2023; Lucey et al., 2023; Sanders & Matsunaga, 2023; Zhang et al., 2023). A supervised machine learning technique trained on nearly 14,000 WDs with Sloan Digital Sky Survey (SDSS) spectra was applied to 100,000 WDs in the Gentile-Fusillo WD catalog with Gaia XP spectra (Vincent et al., 2023). Supervised machine learning requires the use of a labeled dataset where each input feature corresponds to an output label. The dataset trains the machine learning algorithm to recognize patterns and learn the expected outcome for a given input. Once the algorithm is trained, it can be fed unlabeled data for which it will predict the output. In Vincent et al. (2023), the SDSS training dataset included six spectral types for which to classify the WD Gaia XP spectra: DA, DB, DC, DO, DQ, or DZ. Each WD is then assigned a probability for each spectral type. A disadvantage of this technique concerning the DZs is that the supervised machine learning tends to pick out the DZs with the strongest Ca II H&K lines, which are some of the most prominent features in many DZ spectra, thereby excluding or misclassifying DZs with weaker Ca lines. Vincent et al. (2023) quotes a 60% precision rate for the classification of DZs.
We present a different, but complementary, method that utilizes the Gaia XP spectra and an unsupervised machine learning tool called UMAP  [^3] https://umap-learn.readthedocs.io/en/latest/ (McInnes et al., 2020) to organize our sample of 96,134 WDs into identifiable spectral groups. Using unsupervised machine learning to categorize low-resolution objects has already been shown to be successful. A similar manifold learning technique called t-Distributed Stochastic Neighbor Embedding (t-SNE) was implemented for finding metal-poor stars with low-resolution (Rsimilar-$\sim$750) spectra in the Hobby Eberly Telescope Dark Energy Experiment survey (Hawkins et al., 2021). Moreover, the UMAP method in conjunction with Gaia XP spectra was previously utilized by Sanders & Matsunaga (2023) to separate C- and O-rich asymptotic giant branch stars. Unlike the supervised machine learning method (Vincent et al., 2023), unsupervised machine learning methods do not require a prior training dataset. Instead, they are free to interpret patterns and structures in the input features without instruction on the expected outcomes. Unsupervised machine learning is especially useful for detection of anomalous data points and clustering of data with similar attributes.
@@ 4.1 Cool DZs
Figure 3: The UMAP space with all DAs, DBs, DOs, and DZs from the MWDD plotted with blue circles, gold triangles, purple stars, and red diamonds, respectively.
@@
There are approximately 465 total candidates within this cool DZ region with the potential for harboring multiple metal species. Currently, there are a few dozen known DZs with five or more metal species present in their atmospheres (Bonsor et al., 2023; Hollands et al., 2017, 2018). 375 of the 465 we have identified in the cool DZ island are not accounted for in MWDD, meaning that we could potentially increase the number of known heavily polluted WDs tenfold. About 80% of the 375 are also classified as high-probability ($P>$ 65%) DZs in Vincent et al. (2023). The other 20% show significantly less Ca absorption in their XP spectra compared to the high-probability DZs in the Vincent catalog, but still exhibit metal pollution. So while there is good correlation between the two methods, UMAP still uncovered several potential cool DZs that the supervised machine learning technique overlooked.
@@ 4.3.1 ZZ Cetis
In Fig. 7 we plot a zoom-in of the ZZ Ceti instability strip in the UMAP space with the color gradient according to mass. The masses come from the Gentile-Fusillo catalog and are calculated using mass-radius relations for C/O-core WDs with thick H atmospheres (Bédard et al., 2020). The higher mass WDs tend to clump up near the inner boundary of the DA horseshoe and could potentially reveal many new massive pulsator candidates. Currently, there are 6 known pulsators with masses greater than 1 M (Kanaan et al., 1998; Hermes et al., 2013; Curd et al., 2017; Vincent et al., 2020; Kilic et al., 2023). The UMAP candidate selection method reveals over 20 candidates with masses $\gtrsim$ 1 M and Teff within the temperature range where pulsations should be detectable. All candidates have $V_{\sigma}>2$.
@@ 4.3.2 DB WDs
There are two flavors of He-atmosphere WDs: DOs and DBs. Hot He-rich WDs exhibit singly ionized (one electron is stripped) He II lines and are labeled as DOs whereas He-rich WDs below 40,000 K exhibit neutral He I lines and are labeled as DBs.
@@
One mystery in regard to the DBs is the so-called DB gap. There is a near absence of DBs in the 30,000 K $\leq$ Teff $\leq$ 45,000 K range, postulated to arise from mixing of a thin H layer by the convecting He layer below (Koester & Kepler, 2015). After entering a quiescent stage of evolution, a small amount of H present in a DO atmosphere could float to the surface, making the WD appear as a hot DA. It would later be diluted by the onset of He convection around 30,000 K, therefore explaining the existence of the DB gap. However, because there are still DBs within the gap, albeit a small amount, the settling out of H may not occur in all DOs as some of them may have very little to no H in their atmospheres to begin with (Koester & Kepler, 2015).
@@
There are 3694 WDs within the region outlined in the top panel of Fig. 8, 537 of which are classified as DBs in MWDD and 13 as DAs. This means that there potentially could be 3400 previously undiscovered DB WDs. 132 of the WDs included in the DB tip have temperatures between 30,000 and 40,000 K, according to the He-atmosphere photometric fits from the Gentile-Fusillo catalog. Approximately 10 DBs within the DB gap have been discovered thus far (Eisenstein et al., 2006; Kleinman et al., 2013), but our classification method could potentially increase the number of WDs within the DB gap by more than tenfold. Perhaps this could help lift the veil of the DB gap enigma.

### v2

@@ 1 Introduction
The death of a low-mass ($\lesssim$ 8 M) main-sequence star culminates in the ejection of its outer layers in a planetary nebula and the collapse of its core into a white dwarf (WD). Typical mass WDs are extremely dense and have very high surface gravities (log $g$ $\sim$ 8.0, central density $\approx 10^{6}$ $\mathrm{g/cm^{3}}$), equivalent to about 100,000 times that of Earth. As such, WDs are chemically stratified, meaning lighter elements like H and He rise to the surface and heavier elements (e.g., C, O, Ca, Mg, Fe, etc.) sink to the core. The majority of WDs are expected to consist of carbon-oxygen cores and thin upper layers of He and H that make up only $\sim$ 1% of the total WD mass. For most WD spectra, we expect the presence of H (DA) or He I (DB) absorption lines, or no spectral lines (DC) if the WD is cold enough to no longer excite atoms above their ground state ($\lesssim$ 11,000 K for DBs and $\lesssim$ 5,000 K for DAs). However, some WDs, especially cooler WDs, have been found with absorption features from heavier elements in their atmospheres. This has been interpreted as evidence for evolved planetary systems and surviving minor planets (Debes & Sigurdsson, 2002; Jura, 2003; Zuckerman et al., 2007; Koester et al., 2014).
@@
About 25 – 50% of all WDs below 20,000 K are polluted by heavy metals (Zuckerman et al., 2010; Koester et al., 2014; Badenas-Agusti et al., 2023). Observations of intact planets around WDs (Vanderburg et al., 2020; Mullally et al., 2024) and transiting planetary debris (Vanderburg et al., 2015; Vanderbosch et al., 2020, 2021) further lends the notion that planetary systems can persist through to the WD phase. Planets and planetesimals that survive the late-stage evolution of their host stars and enter the WD’s Roche radius, possibly through orbital disruptions during mass loss in the asymptotic giant phase and dynamical interactions with other planets (Maldonado et al., 2020), become tidally disrupted and the debris is subsequently accreted onto the WD. Since the discovery of the first polluted WD, nearly 1400 WDs have been spectrally classified as metal-polluted (DZ, or DAZ, DZA, DBZ, DZB, etc. depending on the predominant lines in their spectra) (Dufour et al., 2017), with a majority only exhibiting Ca pollution.
@@
The most recent Gaia data release (DR3) in 2022 (Gaia Collaboration et al., 2023) revealed nearly 1.5 billion sources with full astrometry, an increase of nearly 136 million sources from the prior data release in 2019. The plethora of data from Gaia has enabled advances in areas such as Galactic archaeology (Poggio et al., 2018; Lucey et al., 2023; Deason & Belokurov, 2024), stellar evolution (Jao et al., 2018; Fouesneau et al., 2023), and measuring cosmic distances (Bailer-Jones et al., 2021; Riess et al., 2021; Ripepi et al., 2023). Gaia has also been pivotal in the field of WD astrophysics, increasing the number of WDs by nearly 100,000 after the second data release (Gentile Fusillo et al., 2019), which contributed to such advancements as the discovery of the WD crystallization branch and cooling anomaly (Tremblay et al., 2019; Cheng et al., 2019) and the development of a technique for finding new variable WDs using Gaia excess photometric scatter (Guidry et al., 2021). The Gaia observations consist of broadband (G), blue passband (BP), and red passband (RP) photometry that cover the wavelength ranges 330–1050 nm, 330–680 nm, and 640–1050 nm, respectively. In earlier data releases, BP and RP were simple photometric magnitudes. However, Gaia DR3 integrated low resolution ($R=\lambda/\Delta\lambda\sim$ 70) blue and red grism spectra to create the BP and RP photometry, and released these spectra for 220 million sources. Despite the low resolution of the BP and RP (XP) spectra, they are invaluable for characterizing atmospheric parameters like effective temperature (Teff), surface gravity (log $g$), metallicity, and line identification for vast quantities of stars (Andrae et al., 2023; Zhang et al., 2023; Vincent et al., 2024).
@@ 3.1 Categorizing Gaia DR3 WDs with UMAP
It has already been shown that the Gaia XP spectra can be exploited through machine learning to label and infer stellar parameters for vast quantities of stars (Andrae et al., 2023; Lucey et al., 2023; Sanders & Matsunaga, 2023; Zhang et al., 2023). A supervised machine learning technique trained on nearly 14,000 WDs with Sloan Digital Sky Survey (SDSS) spectra was applied to 100,000 WDs in the Gentile-Fusillo WD catalog with Gaia XP spectra (Vincent et al., 2024). Supervised machine learning requires the use of a labeled dataset where each input feature corresponds to an output label. The dataset trains the machine learning algorithm to recognize patterns and learn the expected outcome for a given input. Once the algorithm is trained, it can be fed unlabeled data for which it will predict the output. In Vincent et al. (2024), the SDSS training dataset included six spectral types for which to classify the WD Gaia XP spectra: DA, DB, DC, DO, DQ, or DZ. Each WD is then assigned a probability for each spectral type. A disadvantage of this technique concerning the DZs is that the supervised machine learning tends to pick out the DZs with the strongest Ca II H&K lines, which are some of the most prominent features in many DZ spectra, thereby excluding or misclassifying DZs with weaker Ca lines. Vincent et al. (2024) quotes 60% precision, or true positive to false positive + true positive ratio, but $>$ 85% recall, or true positive to true positive + false negative ratio, for the classification of DZs. Another supervised machine learning technique using Random Forest classification applied to Gaia WDs within 100 pc with XP spectra (García-Zamora et al., 2023) yielded 90% precision for DZs, but a lower recall of $<$ 50%. They assigned spectral types to 9,446 previously unclassified WDs, 132 of which were classified as DZ, using WDs in MWDD as a training set.
We present a different, but complementary, method that utilizes the Gaia XP spectra and an unsupervised machine learning tool called UMAP  [^3] https://umap-learn.readthedocs.io/en/latest/ (McInnes et al., 2020) to organize our sample of 96,134 WDs into identifiable spectral groups. Using unsupervised machine learning to categorize low-resolution objects has already been shown to be successful. A similar manifold learning technique called t-Distributed Stochastic Neighbor Embedding (t-SNE) was implemented for finding metal-poor stars with low-resolution (Rsimilar-$\sim$750) spectra in the Hobby Eberly Telescope Dark Energy Experiment survey (Hawkins et al., 2021). Moreover, the UMAP method in conjunction with Gaia XP spectra was previously utilized by Sanders & Matsunaga (2023) to separate C- and O-rich asymptotic giant branch stars. Unlike the supervised machine learning method (Vincent et al., 2024), unsupervised machine learning methods do not require a prior training dataset. Instead, they are free to interpret patterns and structures in the input features without instruction on the expected outcomes. Unsupervised machine learning is especially useful for detection of anomalous data points and clustering of data with similar attributes.
@@ 4.1 Cool DZs
Figure 3: The UMAP space with all DAs, DBs, DOs, and DZs from the MWDD overplotted with blue circles, gold triangles, purple stars, and red diamonds, respectively.
@@
There are approximately 465 total candidates within this cool DZ region with the potential for harboring multiple metal species. Currently, there are a few dozen known DZs with five or more metal species present in their atmospheres (Bonsor et al., 2023; Hollands et al., 2017, 2018). 375 of the 465 we have identified in the cool DZ island are not accounted for in MWDD, meaning that we could potentially increase the number of known heavily polluted WDs tenfold. About 80% of the 375 are also classified as high-probability ($P>$ 65%) DZs in Vincent et al. (2024). The other 20% show significantly less Ca absorption in their XP spectra compared to the high-probability DZs in the Vincent catalog, but still exhibit metal pollution. Similarly, 104 WDs found with the Random Forest method (García-Zamora et al., 2023) show up in our polluted region, but only 75% are classified as DZ. The other 25% classified as DC or DA also exhibit metal pollution in their XP spectra, but are less dominated by Ca absorption. So while there is good correlation between the unsupervised and supervised methods, UMAP still uncovered several potential cool DZs that the supervised machine learning techniques overlooked.
@@ 4.3.1 ZZ Cetis
In Fig. 7 we plot a zoom-in of the ZZ Ceti instability strip in the UMAP space with the color gradient according to mass. The masses come from the Gentile-Fusillo catalog and are calculated using mass-radius relations for C/O-core WDs with thick H atmospheres (Bédard et al., 2020). The higher mass WDs tend to clump up near the inner boundary of the DA horseshoe and could potentially reveal many new massive pulsator candidates. Currently, there are at least 6 known pulsators with masses greater than 1 M (Kanaan et al., 1998; Hermes et al., 2013; Curd et al., 2017; Vincent et al., 2020; Kilic et al., 2023). The UMAP candidate selection method reveals over 20 candidates with masses $\gtrsim$ 1 M and Teff within the temperature range where pulsations should be detectable. All candidates have $V_{\sigma}>2$.
@@ 4.3.2 DB WDs
DOs and DBs are He-atmosphere WDs that feature two different types of He absorption. Hot He-rich WDs exhibit singly ionized (one electron is stripped) He II lines and are labeled as DOs whereas He-rich WDs below 40,000 K exhibit neutral He I lines and are labeled as DBs.
@@
One mystery in regard to the He-atmosphere WDs is the so-called DB gap, marked by a near absence of DBs in the 30,000 K $\leq$ Teff $\leq$ 45,000 K range (Koester & Kepler, 2015; Torres et al., 2023). After entering a quiescent stage of evolution, a small amount of H present in a DO atmosphere could float to the surface, making the WD appear as a hot DA. It would later be diluted by the onset of He convection around 30,000 K, therefore explaining the existence of the DB gap. However, because there are still DBs within the gap, albeit a small amount, the settling out of H may not occur in all DOs as some of them may have very little to no H in their atmospheres to begin with (Koester & Kepler, 2015).
@@
There are 3694 WDs within the region outlined in the top panel of Fig. 8, 537 of which are classified as DBs in MWDD and 13 as DAs. This means that there potentially could be 3400 previously undiscovered DB WDs. 132 of the WDs included in the DB tip have temperatures between 30,000 and 40,000 K, according to the He-atmosphere photometric fits from the Gentile-Fusillo catalog. Approximately 70 DBs and DBAs (mixed He and H atmospheres) within the DB gap have been discovered thus far (Eisenstein et al., 2006; Kleinman et al., 2013), but our classification method could potentially double the number of WDs within this temperature range.
@@ 7 Acknowledgements
KH acknowledges Vasily Belokurov, JJ Hermes, and Nicola Gentile Fusillo for insightful conversations that led to this work. KH acknowledges support from the National Science Foundation grant AST-2108736. This work was performed in part at the Simons Foundation Flatiron Institute’s Center for Computational Astrophysics during KH’s tenure as an IDEA Fellow.